# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ Abstract
We present a decision-theoretic justification for viewing the question of how to best choose where to experiment in order to optimize external validity as a $k$-median (clustering) problem, a popular problem in computer science and operations research. We present conditions under which minimizing the worst-case, welfare-based regret among all nonrandom schemes that select $k$ sites to experiment is approximately equal—and sometimes exactly equal—to finding the $k$ most central vectors of baseline site-level covariates. The $k$-median problem can be formulated as a linear integer program. Two empirical applications illustrate the theoretical and computational benefits of the suggested procedure.
@@ 1 Introduction
A common concern in randomized evaluations of new policies is their external validity; that is, whether the estimated effects of a policy intervention carry over to new samples or populations (Vivalt, 2020; Gechter, 2024; Duflo et al., 2007, Chapter 8). There is a large body of literature arguing that the external validity of randomized evaluations can be improved by explicitly incorporating this goal into the experimental design by, for example, carefully deciding where to experiment; see Degtiar and Rose (2023) for an overview and references, and also the recent work of Chassang and Kapon (2022). For instance, if a researcher has access to multiple sites to experimentally evaluate a new policy, it is possible to use the information available before the evaluation (such as site-level characteristics) to nonrandomly select one site—or more—that are “representative” of the populations of interest. Recently, both Egami and Lee (2024) and Gechter et al. (2024) have shown that a research design that nonrandomly selects experimental sites—referred to broadly in the literature as purposive sampling (Cook et al., 2002, p. 511)—is useful to improve the external validity of randomized evaluations.
@@
This paper presents a theoretical justification for viewing the question of how to best design a purposive sampling scheme—or, equivalently, how to best nonrandomly select $k$ sites to experiment—with the goal of optimizing external validity as a $k$-median problem. This is a classical problem in computer science and operations research (Williamson and Shmoys, 2011; Cohen-Addad et al., 2022). Broadly speaking, in a $k$-median problem, there is a set of facilities and a set of clients; the goal is to open at most $k$ facilities and connect each client to at least one facility at minimal total connection cost. We next provide more specific details on how these problems relate.
@@
Statistical Decision Theory: Under conditions that will be explained clearly in Sections 2 and 3, Theorem 1 in this paper shows that the worst-case (welfare-based) regret of any purposive sampling scheme that selects at most $k$ sites to experiment is approximately equal—and sometimes exactly equal—to the objective function of a $k$-median problem with the following features. First, the sites available for experimentation are treated as facilities. Second, the sites where the policy maker would like to implement the new policy are treated as clients. Third, the connection cost between clients and facilities is proportional to the Euclidean distance between the available site-level covariates. Importantly, we show that the solution of the $k$-median problem is exactly “minimax-regret” optimal (among all purposive sampling schemes) when i) the candidate sites for experimentation and the policy-relevant sites are disjoint (Remark 4) and ii) the treatment effect heterogeneity across sites accommodated in the parameter space is substantial (in a sense Theorem 1 makes precise). When ii) holds but i) does not, we show that the optimized value of the $k$-median problem only approximates the minimax-regret value, and that the approximation error improves as either the number of policy-relevant sites increases or the experiments conducted become more precise (Remark 5). The link with the $k$-median problem established in this paper thus shows that selecting the $k$ sites that have the most central vector of covariates tends to optimize external validity (in a minimax-regret sense).
@@
In order to formalize this connection, we leverage recent developments in the literature of treatment choice problems with partial identification; for example Yata (2021); Ishihara and Kitagawa (2021); Montiel Olea et al. (2023). Note first that in the $k$-median problem, each client is typically connected to only one facility (otherwise, the connection cost would not be minimized). In the context of the site selection problem, a connection between a site $i$ (where an experiment was conducted) and a site $j$ (where no experimentation occurred) means that the estimated effects obtained in site $i$ are used to inform policy decisions in site $j$. This means that even when experimental outcomes in $k>1$ sites are available, the solution to the $k$-median problem would typically prescribe the policy maker to only use the information from the site with the smallest connection cost: the nearest neighbor. But when is it decision-theoretically optimal for a policy maker to behave in this way? After $k$ sites have been selected for experimentation, the policy maker faces the “evidence aggregation” problem introduced in Ishihara and Kitagawa (2021) and recently discussed in Yata (2021), Christensen et al. (2022) and Montiel Olea et al. (2023). That is, the policy maker must decide how to use the estimated treatment effects in different—and potentially heterogeneous—sites to make a treatment choice (implement the new policy or preserve the status quo) in the site of interest. It is known that there are conditions under which it is minimax-regret optimal for the policy maker to base decisions on the nearest neighbor (the site with the most similar site-level characteristics), provided the true treatment effects are allowed to vary substantially as a function of site-level covariates; see, for example, Proposition 1 in Montiel Olea et al. (2023). Moreover, in this case, the optimized worst-case regret is proportional to the distance between the baseline covariates of the site of interest and those of its nearest neighbor. Thus, these results on treatment choice problems with partial identification are fundamental to provide a decision-theoretic justification for the use of the $k$-median (clustering) problem to optimize external validity.
@@
Algorithms: The connection with the $k$-median problem clarifies the problem’s difficulty but also suggests efficient algorithms. To see the need for those, recall that any purposive sampling scheme optimizes over “$n$ choose $k$” potential site combinations, where $n$ is the total number of sites available for experimentation. Optimization over purposive sampling schemes also requires the evaluation of some measure of performance that depends on the dimension of the site-level covariates, $d$. Thus, optimally choosing a purposive sampling scheme by simply evaluating the performance of each combination is costly when “$n$ choose $k$” or $d$ is large.
@@
Conceptually, the connection with the $k$-median problem allows us to understand the computational complexity of finding a minimax-regret optimal purposive sampling scheme under the conditions of Theorem 1. Since the $k$-median problem is known to be NP-hard (Kariv and Hakimi, 1979; Megiddo and Supowit, 1984; Cohen-Addad et al., 2018), there is no algorithm whose computational time scales polynomially in the problem’s inputs $(d,k,n)$.
@@
However, from a practical perspective, it is also known that the $k$-median problem admits a linear integer program formulation (Williamson and Shmoys, 2011, Chapter 7.7, p. 185), and is routinely solved with off-the-shelf algorithms such as different versions of the branch-and-bound method (Bertsimas and Weismantel, 2005, Chapter 11.1). In addition, different branch-and-bound algorithms either find a solution with provable optimality or, if stopped early, generate a report on the suboptimality of the solution found (Bertsimas et al., 2016). As we explain later, in our application with $n=41$, $d=13$, any problem for $k\in\{1,\ldots,10\}$ can be solved to provable optimality in just a few seconds using a personal laptop (see Figure 5 in Section 5).
@@
Related Literature: Our results build on two recent papers that presented novel purposive sampling strategies to select experimental sites so as to optimize external validity. Gechter et al. (2024) present an elegant decision-theoretic approach that frames external validity as a policy problem and—under the assumption that the policy maker has a priori information about the effects of the new policy across sites—recommend a Bayesian approach for choosing where to experiment. Egami and Lee (2024) use the principle behind synthetic control (Abadie et al., 2010) to recommend what they call the synthetic purposive sampling of sites; specifically, they select good “donors,” defined as sites whose weighted average of observed baseline characteristics are close to the characteristics of the sites of interest. It is important to note that our decision-theoretic framework leans heavily on Gechter et al. (2024), but also that the sites selected through the $k$-median problem can be interpreted as a degenerate synthetic purposive sampling strategy, whereby each unit’s associated synthetic unit is just its nearest neighbor.
@@
This paper also contributes to the literature arguing that a “modern, decision-theoretic framework can help clarify important practical questions of experimental design” (Banerjee et al., 2017). Although decision-theoretic approaches to external validity are recent, a large body of work used statistical decision theory to analyze other aspects of experimental design, such as sample size determination (Raiffa and Schlaifer, 1961; Manski and Tetenov, 2016, 2019; Azevedo et al., 2020, 2023; Hu et al., 2024). Finally, our notion of external validity is conceptually related to several areas of research in econometrics, machine learning, and statistics such as domain adaptation (Mansour et al., 2009; Ben-David et al., 2010), distributional shifts (Duchi and Namkoong, 2021; Sugiyama et al., 2007; Adjaho and Christensen, 2022), learning under biased sampling (Sahoo et al., 2022), and cross-domain transfer estimation and performance (Andrews et al., 2022; Menzel, 2023). To the best of our knowledge, none of these papers contain decision-theoretic analyses of where to experiment.
@@
Outline: This paper is organized as follows. Section 2 introduces the formal framework. Section 3 presents the main result of this paper (Theorem 1) and explains how the $k$-median problem arises naturally when trying to incorporate external validity concerns into the decision problem of where to experiment. Section 4 presents the linear integer program formulation of the $k$-median problem. Section 5 presents two illustrative empirical applications. Section 6 considers extensions of the baseline model. Section 7 concludes. Proofs of the main results can be found in Appendix A. Additional results are collected in the Supplementary Appendix.
@@ 2 Setting up the Decision Problem
2.1 Notation
A policy maker considers a set of $S\in\mathbb{N}$ candidate sites to evaluate and, eventually, implement a new policy of interest. The sites are indexed by $s\in\mathcal{S}\equiv\{1,\ldots,S\}$. In order to accommodate situations in which the policy maker is not necessarily able to experiment in all the candidate sites, we assume there is a nonempty subset $\mathcal{S}_{E}\subseteq\mathcal{S}$ of what we term experimental sites. [^2] As discussed in Allcott (2015), there are often systematic reasons determining the eligibility of certain sites for experimentation. For example, in microfinance RCT studies, experiments often require large sample sizes and well-managed microfinance institutions (MFIs), characteristics more commonly found in older and larger institutions. To qualify for clinical trials involving a new surgical procedure, hospitals and surgeons need to have both experience in the procedure and a history of low mortality rates. Throughout the paper, and to avoid a trivial instance of the site selection problem, we assume that there are at least two experimental sites (i.e., $\textrm{card}(\mathcal{S}_{E})\geq 2$). It is also possible that institutional restrictions preclude the eventual implementation of the policy of interest in all of the candidate sites. Thus, it will be convenient to denote by $\mathcal{S}_{P}\subseteq\mathcal{S}$ the nonempty set of policy or policy-relevant sites.
In principle—and in order to accommodate different scenarios that could arise in empirical work—we allow for the case in which there is overlap between experimental and policy sites (i.e., $\mathcal{S}_{E}\cap\mathcal{S}_{P}\neq\emptyset$). We note, however, that the case in which these two sets are disjoint (i.e., $\mathcal{S}_{E}\cap\mathcal{S}_{P}=\emptyset$) will be of particular interest. This is because when there is no overlap between experimental and policy sites, the extrapolation problem faced by the policy maker becomes evident: a policy decision needs to be made in sites where there has never been a previous experimental evaluation of the policy of interest.
For each site, the policy maker observes a vector of site characteristics $X_{s}\in\mathbb{R}^{d}$ that may affect the treatment effect. Thus, we allow for treatment effect heterogeneity across sites, but we restrict this heterogeneity by assuming that it depends on observable characteristics. Specifically, let the function $\tau:\mathbb{R}^{d}\rightarrow\mathbb{R}$ define conditional (on $X$) average treatment effects. We posit that any pair of sites with “similar” observed characteristics also have “similar” treatment effects, formally by assuming that $\tau$ is a Lipschitz function (with respect to the Euclidean norm) with known constant $C$. We formally state this assumption in Section 2.4. It is not innocuous, but we will argue that it can be replaced by other continuity-like conditions (such as Hölder continuity) while maintaining the main message of the paper. [^3] In Section 6.3, we discuss how our results extend to other functional classes that could be used to restrict treatment effect heterogeneity, such as collections of functions that belong to a convex and centrosymmetric space. Such restrictions were recently used in the econometrics literature to analyze estimation, inference, and other decision problems that arise in a nonparametric regression setup (Yata, 2021; Armstrong and Kolesár, 2018). We let $\textrm{Lip}_{C}(\mathbb{R}^{d})$ denote the space of all Lipschitz functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with constant $C$.

### v2

@@ Abstract
We present a decision-theoretic justification for viewing the question of how to best choose where to experiment in order to optimize external validity as a $k$-median problem, a popular problem in computer science and operations research. We present conditions under which minimizing the worst-case, welfare-based regret among all nonrandom schemes that select $k$ sites to experiment is approximately equal—and sometimes exactly equal—to finding the $k$ most central vectors of baseline site-level covariates. The $k$-median problem can be formulated as a linear integer program. Two empirical applications illustrate the theoretical and computational benefits of the suggested procedure.
@@ 1 Introduction
A common concern in randomized evaluations of new policies is their external validity; that is, whether the estimated effects of a policy intervention carry over to new samples or populations (Chapter 8, Duflo et al., 2007; Al-Ubaydli and List, 2012; List, 2020; Vivalt, 2020; Gechter, 2024). A recent literature argues that the external validity of randomized evaluations can be improved by explicitly incorporating this goal into the experimental design by, for example, carefully deciding where to experiment; see Degtiar and Rose (2023) for an overview and references and Chassang and Kapon (2022) for some recent work. For instance, if a researcher has access to multiple sites to experimentally evaluate a new policy, it is possible to use the information available before the evaluation (such as site-level characteristics) to nonrandomly select one or more sites. Recently, Egami and Lee (2024) and Gechter et al. (2024) argued that a research design which nonrandomly selects experimental sites—referred to broadly in the literature as purposive sampling (Cook et al., 2002, p. 511)—may improve the external validity of randomized evaluations.
@@
We present a decision-theoretic justification for viewing the question of how to best nonrandomly select $k$ sites to experiment—or, equivalently, how to best design a purposive sampling scheme—with the goal of optimizing external validity as a $k$-median problem. This is a classical problem in computer science and operations research (Williamson and Shmoys, 2011; Cohen-Addad et al., 2022). Broadly speaking, in a $k$-median problem, there is a set of facilities and a set of clients; the goal is to open at most $k$ facilities and connect each client to at least one facility at minimal total connection cost. The $k$-facilities that solve the $k$-median problem are the locations that are closest to the clients, where closeness is defined by the total connection cost. As we will explain below, the $k$ sites selected for experimentation by solving the $k$-median problem will be the $k$ sites that are, broadly speaking, the most representative of the target populations to which the new policy will be scaled up. We next provide more specific details on how we use statistical decision theory to establish a formal connection between the site-selection problem to optimize external validity and the k-median problem.
The Site-Selection Problem: Suppose that some experimental sites can be used by a policymaker (or a researcher) to gather experimental evidence on the effectiveness of a policy or treatment of interest. Because of logistical or budget constraints, we assume that the policymaker can choose at most $k$ sites to experiment. Based on the experimental outcomes of these sites, the policymaker will decide whether it is worthwhile to scale up the policy or treatment of interest in a set of policy-relevant sites. The experimental and policy-relevant sites may or may not overlap. How should the policymaker decide where to experiment? Gechter et al. (2024) formalize the site-selection problem using a statistical decision theory framework in which, after experimentation, the policymaker observes a vector of unbiased and normally distributed estimates of the policy effects of interest at the selected experimental sites. The policy effects are allowed to be heterogeneous across sites. The policymaker is assumed to have access to a vector of baseline site-level covariates for all sites. [^2] The baseline model assumes that the policymaker has no access to individual-level covariates. This means that the policymaker does not have enough information to estimate conditional average treatment effects (CATEs) that could be transported (e.g., via reweighting) to learn the treatment effects of interest at policy-relevant sites. At the end of Section 2.2, we discuss how individual-level covarites, if available, could be used in the site-selection problem.
@@
Connection to the k-median Problem: Under conditions that will be explained clearly in Sections 2 and 3, our main result (Theorem 1) shows that the worst-case (welfare-based) regret of any purposive sampling scheme that experiments in at most $k$ sites is approximately equal—and sometimes exactly equal—to the objective function of a $k$-median problem with the following features: the sites available for experimentation are treated as facilities; the sites where the policymaker would like to implement the new policy are treated as clients; the connection cost between clients and facilities is proportional to the Euclidean distance between corresponding site-level covariate values. Importantly, the solution of the $k$-median problem is exactly minimax-regret optimal (among purposive sampling schemes) when i) the candidate sites for experimentation and the policy-relevant sites are disjoint and ii) the treatment effect heterogeneity across sites accommodated in the parameter space is substantial (in a sense Theorem 1 makes precise). When ii) holds but i) does not, we show that the optimized value of the $k$-median problem only approximates the minimax-regret value, and that the approximation error improves as either the number of policy-relevant sites increases or the experiments conducted become more precise. The link with the $k$-median problem established in this paper thus shows that selecting the $k$ sites that have the most central vector of covariates tends to optimize external validity (in a minimax-regret sense).
@@
In order to formalize the connection between the site selection problem and the $k$-median problem, we leverage recent developments in the literature on treatment choice problems with partial identification (Yata, 2021; Ishihara and Kitagawa, 2021; Montiel Olea et al., 2025). Note first that in the $k$-median problem, each client is typically connected to only one facility (otherwise, the connection cost would not be minimized). In the context of the site selection problem, a connection between a site $i$ (where an experiment was conducted) and a site $j$ (where no experimentation occurred) means that the estimated effects obtained in site $i$ are used to inform policy decisions in site $j$. This means that even when experimental outcomes in $k>1$ sites are available, the solution to the $k$-median problem would prescribe the policymaker to only use the information from the site with the smallest connection cost: the nearest neighbor. But when is it decision-theoretically optimal for a policymaker to behave in this way? After $k$ sites have been selected for experimentation, It turns out that, conditionally on having selected experimental sites, the policymaker faces the “evidence aggregation” problem introduced in Ishihara and Kitagawa (2021) and recently discussed in Yata (2021), Christensen et al. (2022), and Montiel Olea et al. (2025). This literature has established conditions under which it is minimax-regret optimal to base decisions on the nearest neighbor’s data, provided the true treatment effects are allowed to vary substantially as a function of site-level covariates; see, for example, Montiel Olea et al. (2025, Proposition 1). Moreover, in this case, the optimized worst-case regret is proportional to the distance between the baseline covariates of the site of interest and those of its nearest neighbor. Thus, these results on treatment choice problems with partial identification are a building block of our decision-theoretic justification for the use of the $k$-median (clustering) problem to optimize external validity.
@@
Algorithms: The connection with the $k$-median problem clarifies the problem’s difficulty but also suggests efficient algorithms. To see the need for those, recall that any purposive sampling scheme optimizes over “$n$ choose kk” potential site combinations, where $n$ is the total number of sites available for experimentation. Optimization over purposive sampling schemes also requires the evaluation of some measure of performance that depends on the dimension of the site-level covariates, $d$. Thus, optimally choosing a purposive sampling scheme by simply evaluating the performance of each combination is costly when “$n$ choose kk” or $d$ is large.
@@
Conceptually, the connection with the $k$-median problem allows us to understand the computational complexity of finding a minimax-regret optimal purposive sampling scheme under the conditions of Theorem 1. Since the $k$-median problem is known to be NP-hard (Kariv and Hakimi, 1979; Megiddo and Supowit, 1984; Cohen-Addad et al., 2018), there is no algorithm for finding a minimax-regret optimal purposive sampling scheme whose computational time scales polynomially in all of the problem’s inputs; namely, $(d,k,n)$.
@@
However, from a practical perspective, it is known that the $k$-median problem admits a linear integer program formulation (Williamson and Shmoys, 2011, Chapter 7.7, p. 185), and is routinely solved using off-the-shelf algorithms such as different versions of the branch-and-bound method (Bertsimas and Weismantel, 2005, Chapter 11.1). In addition, different branch-and-bound algorithms either find a solution with provable optimality or, if stopped early, generate a report on the suboptimality of the solution found (Bertsimas et al., 2016). As we explain later, in our application with $n=41$, $d=13$, any problem for $k\in\{1,\ldots,10\}$ can be solved to provable optimality in just a few seconds using a personal laptop (see Figure 5 in Section 5).
@@
Related Literature: Our results build on two recent papers that present novel purposive sampling strategies to select experimental sites so as to optimize external validity. Gechter et al. (2024) present an elegant decision-theoretic approach that frames external validity as a policy problem and—under the assumption that the policymaker has a priori information about the effects of the new policy across sites—recommend a Bayesian approach for choosing where to experiment. Egami and Lee (2024) use the principle behind synthetic control (Abadie et al., 2010) to recommend the synthetic purposive sampling of sites; specifically, they select good donor sites whose weighted average of covariates is close to those of the sites of interest. [^3] The idea of using the synthetic control method for experimental design was first introduced in Abadie and Zhao (2021). It is important to note that the site selection achieved by solving the $k$-median problem can be interpreted as a degenerate synthetic purposive sampling strategy, whereby each unit’s associated synthetic unit is just its nearest neighbor.
@@
We also contribute to the literature arguing that a “modern, decision-theoretic framework can help clarify important practical questions of experimental design” (Banerjee et al., 2017). Although decision-theoretic approaches to external validity are recent, a large body of work used statistical decision theory to analyze other aspects of experimental design such as sample size determination (Raiffa and Schlaifer, 1961; Manski and Tetenov, 2016, 2019; Azevedo et al., 2020, 2023; Hu et al., 2024). Finally, our notion of external validity is conceptually related to areas of research in econometrics, machine learning, and statistics such as domain adaptation (Mansour et al., 2009; Ben-David et al., 2010), distributional shifts (Duchi and Namkoong, 2021; Sugiyama et al., 2007; Adjaho and Christensen, 2022), learning under biased sampling (Sahoo et al., 2022), and cross-domain transfer estimation and performance (Andrews et al., 2022; Menzel, 2023). To the best of our knowledge, none of these papers contains decision-theoretic analyses of where to experiment. Our site selection problem is also related to optimal regression design (e.g., Sacks and Ylvisaker 1984) and kriging (e.g., Stein 1999); both consider a mean square error criterion with either Bayes or minimax optimality, while our approach focuses on minimax welfare regret optimality. See also Karmakar (2022) and references therein for the study of blocked randomization designs in stratified experiments to improve precision of treatment effect estimates.
@@
Outline: This paper is organized as follows. Section 2 introduces the formal framework. Section 3 presents our main result linking the $k$-median problem to considerations of external validity. Section 4 presents the linear integer program formulation of the $k$-median problem. Section 5 presents two illustrative empirical applications. Section 6 considers extensions of the baseline model. Section 7 concludes. Proofs of the main results can be found in Appendix A. Additional results are collected in the Supplementary Appendix.
@@ 2 Setting up the Decision Problem
2.1 Notation and Assumptions
[…]
We allow for overlap between experimental and policy sites, i.e., $\mathcal{S}_{E}\cap\mathcal{S}_{P}\neq\emptyset$. However, the case in which these sets are disjoint allows for particularly succinct analysis. This case also makes the extrapolation problem particularly stark: A policy decision must be made in sites where no experimental evaluations are available.
For each site, the policymaker observes a vector of site characteristics $X_{s}\in\mathbb{R}^{d}$ that may affect the treatment effect. Thus, we allow for treatment effect heterogeneity across sites, but we restrict this heterogeneity by assuming that it depends on observable site-level characteristics. Specifically, let the function $\tau:\mathbb{R}^{d}\rightarrow\mathbb{R}$ define conditional (on $X$) average treatment effects. We posit that any pair of sites with similar observed characteristics also have similar treatment effects, formally by assuming the following:
Assumption 1.
$\tau$ is a Lipschitz function (with respect to the Euclidean norm) with known constant $C$. That is, for any $x,x^{\prime}\in\mathbb{R}^{d}$, $|\tau(x)-\tau(x^{\prime})|\leq C\|x-x^{\prime}\|,$ where $\|\cdot\|$ denotes the Euclidean norm.
This assumption is not innocuous, but we will argue that it can be replaced by other continuity-like conditions. For example, in Section 6.3.1, we give a version of our main results using a weaker version that includes Hölder continuous functions. Moreover, in Section 6.3.4, we discuss how to further relax this assumption by assuming that $\tau$ belongs to a convex and centrosymmetric space of functions. [^5] Such a restriction has been used recently in the econometrics literature to analyze estimation, inference, and other decision problems that arise in a nonparametric regression setup (Yata, 2021; Armstrong and Kolesár, 2018). In Section 6.3.2, we also discuss how Assumption 1 can be modified to accommodate some forms of unobserved treatment heterogeneity. Perhaps it is important to note at this point that the purpose of our paper is not to present a solution to the site selection problem under the most general set of conditions on the parameter space and the statistical model. Instead, our goal is to show that under reasonable assumptions, the principle of selecting the sites that are most representative of the policy-relevant sites can be fully rationalized using a statistical decision-theoretic framework, and in addition can be implemented using off-the-shelf algorithms in optimization. We let $\textrm{Lip}_{C}(\mathbb{R}^{d})$ denote the space of all Lipschitz functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with constant $C$.
In addition to Assumption 1, we also impose a regularity condition on site-level covariates. That is, we assume that all observed covariates are different:
Assumption 2.
$X_{s}\neq X_{s^{\prime}}\ \forall\ s,s^{\prime}\in\mathcal{S}$.
Even if this were not the case in raw data, one would presumably want to induce it by adding site fixed effects.